Dr. Sand borrowed some money for new furniture. She paid $245. simple interest on a 3.5 year loan at 3.5%. Find the principle.
Please help,
Ellen
: Dr. Sand borrowed some money for new furniture. She paid $245. simple interest on a 3.5 year loan at 3.5%. Find the principle.
Please help,
Ellen

Answer by josmiceli(2018)

The formula is
A = P*(1 + tr)
Principle = P
number of years = t
Amount paid on loan after t years = A
Annual interest rate = r
Interest after t years = A - P
------------------
Given is r = .035
A - P = 245
t = 3.5
------------------
P = A - 245
A = P*(1 + tr)
A = (A - 245)*(1 + tr)
A = (A - 245)*(1 + 3.5*.035)
A = (A - 245)*1.1225
A = 1.1225A - 275.0125
.1225A = 275.0125
A = 2245
P = A - 245
P = 2245 - 245
P = 2000
The principle was $2000 answer
plug this back into formula to check
A = P*(1 + tr)
2245 = 2000*(1 + 3.5*.035)
2245 = 2000*1.1225
2245 = 2245
OK
Question 149841: Dr. Sand borrowed some money for new furniture. She paid $245. simple interest on a 3.5 year loan at 3.5%. Find the principle.
Please help,
Ellen
: Dr. Sand borrowed some money for new furniture. She paid $245. simple interest on a 3.5 year loan at 3.5%. Find the principle.
Please help,
Ellen

Answer by scott8148(2719)

for simple interest __ I=P*R*T

dividing by R*T __ I/(R*T)=P __ 245/(.035*3.5)=P __ 2000=P